Karen walsh has twice as much money invested at 5% simple annual interest as she does at 4%. if her yearly income from these two investments is $350, how much does she have invested at each interest rate?
Answer by mananth(16946)

Part I 4.00% per annum------------- x
Part II 5.00% per annum------------y= 2x
Interest----- $350.00

4.00% x + 5.00% * 2 x = $350.00
0.04 x + 0.05 * 2 x = $350.00
0.04 x + 0.1 x = $350.00
0.14 x = $350.00
/ 0.14
x= $2,500.00 --- Part I
2x = $5,000.00 Part II